# Continuous Speech Synthesis using per-token Latent Diffusion

### v1

@@ 4.3 Ablation Study
T2A HiFi Continuous $d=32$ 4.271 1.157 0.545
T2A HiFi Discrete $Q=12$ 4.203 5.461 0.597
Table 2: Discrete vs continuous models with high-fidelity representations
@@
VAE Sample 4.280 0.739 0.539
VAE NoSample 3.468 1.891 0.613
Table 3: Influence of VAE sampling during training
VAE sampling
VAE models allow the sampling of diverse inputs, unlike the discrete codebooks or Mel spectrograms. This ability may improve the robustness of the model, and better account for the mismatch between training and inference (during inference, the model predictions are based on its previous noisy predictions). To check the influence of VAE sampling, we compare two T2A models - one samples from the VAE distribution $x=\mu+\epsilon\cdot\sigma$ and the other always takes the mean $x=\mu$. The results in Table 3 show a large gap in UTMOS and intelligibility indicating that sampling improves synthesized samples. We listened to audio samples from the VAE-NoSample model, and noticed a gradual addition of speaker-inconsistency artifacts throughout the synthesis. We suspect that the addition of VAE-sampling noise during training made it more robust to the mismatch between training and inference. We also hypothesize that high similarity result of VAE-NoSample is caused by the artifacts described above.
5 Discussion
Compressing complex signals such as audio and images often introduces a tradeoff between perception and generation. For tasks involving perception or understanding, compression can lead to information loss, resulting in degraded performance. However, for generation, compression has proven to be highly effective, as the generative model has to learn a lower-dimensional distribution. Multimodal models typically aspire to work with symmetric representations, in which the input and output representations are identical, as commonly done in language models. Developing generative methods that can operate upon less-compressed representations would alleviate the perception-generation models, and improve multimodal models that operate on symmetric representations. While RVQ is a powerful compression mechanism, capable of providing high-fidelity representations, it may lead to noise quantization. Working with continuous representations can be more robust to noise, as continuous models scale the noise according to its magnitude.
Limitations
The diffusion head inference process is slower than the RVQ prediction heads, as it requires an iterative process for the token sampling. Moreover, it does not allow to measure likelihood or confidence, which can be useful for decoding algorithms such as beam search or confidence based unmasking. Optimal balancing of the discrete and continuous losses in the continuous T2A model is not easy to obtain. During training, the gradients of the discrete semantic loss increase, while the gradients of the continuous diffusion loss decrease.
Future work
Follow-up works can extend our work and develop multimodal models that operate on symmetric representations, and are capable of perception and generation. They can also derive a generation stopping condition that does not rely on any discrete representation. Additionally, future works can implement diverse inference strategies that adapt the number of diffusion steps per token (e.g. more diffusion steps in the first tokens), or develop a quality metric for a diffusion process, to allow decoding algorithms such as beam-search to be used during inference.
Appendix A Additional Results
A.1 GIVT
We also attempted to use GIVT (Tschannen et al., 2023) as an alternative approach for continuous audio generation. GIVT models the next token distribution as a Gaussian mixture model (GMM). We trained a GIVT model to a mixture of 16 Gaussians, which predicts the next continuous acoustic frame. We focused on the ability to produce multi-mode distributions. Figure 8(a) plots the entropy of mixture coefficients in the GMM, which drops quickly to zero during training. This might suggest that the ability to produce multimodal probability distributions is not being leveraged frequently.
(a) GIVT GMM entropy drops to 0.
(b) SoundStorm unmasking methods
Figure 8: Additional results
A.2 MaskGIT inference
In S2A-NAR-Cont, the MaskGIT selection of tokens to unmask is done at random, instead of being based on confidence scores. To check the influence of the unmasking approach, we provide three unmasking criteria for our SoundStorm model: highest confidence, random, and lowest confidence. The test was based on GT semantic tokens (to avoid dependency on semantic token prediction), with the default hyperparameters described in Section 4. We first sampled each token using top-k sampling, and defined the token score to be the softmax probability of the sampled token. We then unmasked tokens based on the score, its inverse, or at random. The results in Figure 8(b) suggest that random selection in SoundStorm yields improved performance over confidence-based selection. This resembles the results seen when using greedy sampling, which leads to a sub-optimal result.

### v2

@@ V-C Ablation Study
Continuous (d=32) 4.271 1.157 0.545
Discrete (Q=12) 4.203 5.461 0.597
TABLE II: Discrete vs. continuous SALAD  with high-fidelity representations. Intelligibility (CER) and Similarity (speaker similarity) are defined in Section. V (Metrics).
VI Discussion
Compressing complex signals like audio and images involves a tradeoff between perception and generation. While compression can degrade perception by losing information, it benefits generation by simplifying the learned distribution. Developing generative methods for less-compressed representations could help mitigate this tradeoff. Though RVQ enables high-fidelity compression, it may quantize noise. Using continuous representations can be more robust to noise, as continuous models scale the noise according to its magnitude.
Future work.  Our approach could be extended by developing multimodal models capable of both perception and generation. Additionally, exploring generation stopping conditions that do not rely on discrete representations would be valuable. Adaptive inference strategies, such as adjusting the number of diffusion steps per token (e.g., using more steps for early tokens), could improve efficiency. Finally, developing quality metrics for the diffusion process could enable the use of decoding algorithms like beam search during inference.
Limitations  The diffusion head inference process is slower than the RVQ prediction heads, as it requires an iterative process for the token sampling. Moreover, it does not allow to measure likelihood or confidence, which can be useful for decoding algorithms such as beam search or confidence based unmasking. Optimal balancing of the discrete and continuous losses in SALADis not easy to obtain. During training, the gradients of the discrete semantic loss increase, while the gradients of the continuous diffusion loss decrease.
Ethical Statement  SALAD is capable of zero-shot voice cloning, which presents potential risks, including misuse for voice spoofing. To mitigate these risks, it is essential to implement strict protocols that verify and ensure the speaker’s consent and authorization before their voice is used in any real-world application, especially when dealing with previously unseen speakers. Safeguards must be in place to prevent unauthorized use and protect individuals’ privacy and identity.
@@ VI Discussion
Sample 4.280 0.739 0.539
NoSample 3.468 1.891 0.613
TABLE III: Influence of VAE sampling during training
-0.2cm
AI Tools Usage  AI-based tools may have assisted in code writing and paraphrasing, but all content was thoroughly reviewed by the authors and used only as a support tool.
Acknowledgment
This research work was partially supported by Mobileye Academic Grant Program and by the Israel Innovation Authority, grant number 78563.